# Predicting Census Survey Response Rates with Parsimonious Additive Models and Structured Interactions

### v4

@@ 5.3 Interpreting important features
We now illustrate how our methodology can guide the practitioner in obtaining a set of features and deriving associated actionable insights into the factors contributing towards low response-rates in surveys. An important aspect of our regularized learning framework is that it provides an automated scheme to identify a collection of models, balancing out the complexity of the model and data-fidelity. To see this, Figure 7[left panel] plots the number of main and interaction effects against the associated prediction error for model (12). The plot illustrates that by trading off a little in the predictive performance (in terms of RMSE, which is shown by the red dashed line), we can limit the number of effects at any level, as desired by the practitioner who intends to obtain a more parsimonious representation of the factors impacting survey response rates. Specifically, if we would like to limit the number of main effects to under $20$, Figure 7(b) shows the top $19$ main effects in the order they enter the model along the regularization path. The definitions of the variables in Figure 7(b) are provided in the Supplement. Our investigation reveals that most (though not all) of these variables also appear in the models reported in prior studies by the Census Bureau – see for example Erdman and Bates (2016) and Kulzick et al. (2019). Interestingly, some of the features discovered by our framework can be interpreted in the context of clusters defined in Bates and Mulry (2011) or mindset solutions in Kulzick et al. (2019), as discussed below.
@@
Figure 7: Estimates available from our model ELAAN-I i.e., model (12): $\ell_{0}$-sparse AMs with main effects and interactions. [Left] number of nonzero effects vs the corresponding RMSE. [Right] Variables corresponding to main effects selected by the model along the regularization path. For visualization purposes, we focus on the top 19 main effects.
Using cluster analysis on the 12 covariates from the HTC score, Bates and Mulry (2011) categorize the tracts into eight distinct clusters: All Around Average I and II, Economically Disadvantaged I and II, Ethnic Enclave I and II, Single Unattached Mobiles, and Advantaged Homeowners. From the cluster analysis in Bates and Mulry (2011), we see that the reported clusters have a wide variation in response rate (the response rate is not used to identify the clusters). Each cluster has some key characteristics (pertaining to a subset of variables) that seem to drive the response rate. From the description of these clusters in Bates and Mulry (2011), we see that the covariates automatically discovered by our framework, as listed in Figure 7[right panel], broadly cover the important characteristics upon which the clusters are based. For example, there is high variability across clusters with respect to the occupancy rate and home ownership status, which are captured by covariates “Owner occupied housing units” and “Total occupied units”. Racial diversity also appears to play a role in cluster formation – in particular, the covariates related to ethnic groups characterize the Ethnic Enclave clusters. Similarly, covariates related to poverty, for example, “Persons below poverty line”, lack of education (“No high school diploma”), and lack of health insurance (“No health insurance”), determine the Economically Disadvantaged clusters. Covariates related to high mobility, for example, “Different housing unit in prior year”, high education (“College or higher degree holders”), marital status (“Single person households”), and nonspousal occupants (“Unmarried couples”), characterize the Single Unattached Mobiles cluster.
@@
Some of the covariates in Figure 7[right] can also be interpreted in terms of the mindset solutions in Kulzick et al. (2019). In a mindset solution, survey respondents are categorized using factor analysis into six mindsets based on their predicted self-response patterns and demographic characteristics. These mindsets are determined by the demographic characteristics such as age, race, income, home-ownership, presence of children in the household, marital status, internet use, English proficiency, and country of birth. The six mindsets are: Eager Engagers, Fence Sitters, Individuals with a Confidentiality, Head Nodders, Wary Skeptics and Disconnected Doubters. The covariates can be also interpreted in terms of eight segmented tracts: Responsive Suburbia, Main Street Middle, Country Roads, Downtown Dynamic, Student and Military Communities, Sparse Spaces, Multicultural Mosaic, and Rural Delta and Urban Enclaves – see Kulzick et al. (2019) for additional details.
@@ 5.4 Conclusions
We propose a framework to obtain models with good predictive accuracy while simultaneously delivering parsimonious and interpretable statistical models. We hope that our framework will help practitioners identify some key factors that influence survey response rates. Insights gathered from our models may be used for targeted follow-up surveys and allocation of resources for advertisements. We hope that our models will help inform the goal of achieving improved census coverage in a timely fashion, while optimizing the operational risks and costs. Our framework can complement and potentially offer an alternative to the LRS metric that is currently used in the US Census Bureau ROAM application.
@@ Acknowledgements
The authors would like to thank Nancy Bates, Tommy Wright and Joanna Fane Lineback from the U.S. Census Bureau for helpful discussions. Thanks also to other researchers from the Centers for Statistical Research and Methodology and Behavioral Science Methods (U.S Census Bureau) for helpful feedback in earlier versions of this work. Any opinions and conclusions expressed herein are those of the authors and do not necessarily represent the views of the U.S. Census Bureau. This research was supported in part, by grants from the Office of Naval Research: ONR-N000141812298 (YIP) and the National Science Foundation: NSF-IIS-1718258 awarded to Rahul Mazumder.
@@
We study how our model performs when the error distribution is (i) skewed, (ii) heteroskedastic. For (i), we consider a log-normal distribution for the error: $\epsilon\sim\mathcal{LN}(-1,0.2546)$. For (ii), we consider a normal distribution for the error: $\epsilon\sim\mathcal{N}(0,0.5092g_{2}(x_{5}))$, where the standard deviation is a function of the 5-th covariate in $\mathbf{x}$. We compare average ISE for ELAAN-I with EBM and GAMI-Net in these scenarios in Table S2. We can observe that ELAAN-I consistently outperforms both EBM and GAMI-Net even when the error distributions are not normally distributed. As expected, we observe a small drop in performance in both skewed and heteroskedastic settings compared to the normal distribution setting. We do not observe any degradation in support recovery metric in Table S3 under skewed setting; however, we do observe a drop in support recovery F1-scores in the heteroskedastic setting. For example, the F1-scores for main and interaction effects for ELAAN-I are $96.4\%$ and $81.7\%$ in the heteroskedastic setting for $N_{train}=400$ compared to $97.4\%$ and $87.3\%$ in the normal distribution setting.
@@
For EBM, we tune the number of interactions in the range $[1,100]$. For GAMI-Net, we tuned the model with number of interactions in the set $\{5,50\}$. For GAMI-Net, we used $200$ batch-size, $0.0001$ learning rate, $500$ epochs (for each stage) and $0.0$ loss threshold.

### v5

@@ 5.3 Interpreting important features
We now illustrate how our methodology can guide the practitioner in obtaining a set of features and deriving associated actionable insights into the factors contributing to low response rates in surveys. An important aspect of our regularized learning framework is that it provides an automated scheme to identify a collection of models, balancing out the complexity of the model and data-fidelity. To see this, Figure 7[left panel] plots the number of main and interaction effects against the associated prediction error for model (8). The plot shows that by trading off slight predictive performance (in terms of RMSE, which is shown by the red dashed line), we can limit the number of effects at any level, as desired by the practitioner who seeks a more parsimonious model. Specifically, if we would like to limit the number of main effects to under $20$, Figure 7(b) shows the top $19$ main effects in the order they enter the model along the regularization path. The definitions of the variables in Figure 7(b) are provided in the Supplement (Ibrahim et al., 2024). Our investigation reveals that most (though not all) of these variables also appear in the models reported in prior studies by the Census Bureau – see, for example , Erdman and Bates (2016) and Kulzick et al. (2019). Interestingly, some features discovered by our framework can be interpreted in the context of clusters defined in Bates and Mulry (2011) or mindset solutions in Kulzick et al. (2019), as discussed below.
@@
Figure 7: Estimates available from our model ELAAN-I i.e., model (8): $\ell_{0}$-sparse AMs with main effects and interactions. [Left] number of nonzero effects vs the corresponding RMSE. [Right] Variables are corresponding to the main effects selected by the model along the regularization path. For visualization purposes, we focus on the top 19 main effects.
Using cluster analysis on the 12 covariates from the HTC score, Bates and Mulry (2011) categorize the tracts into eight distinct clusters: All Around Average I and II, Economically Disadvantaged I and II, Ethnic Enclave I and II, Single Unattached Mobiles, and Advantaged Homeowners. From the cluster analysis in Bates and Mulry (2011), we see that the reported clusters have a wide variation in response rate (the response rate is not used to identify the clusters). Each cluster has some key characteristics (pertaining to a subset of variables) that seem to drive the response rate. From the description of these clusters in Bates and Mulry (2011), we see that the covariates automatically discovered by our framework, as listed in Figure 7[right panel], broadly cover the important characteristics upon which the clusters are based. For example, there is high variability across clusters with respect to the occupancy rate and home ownership status, which are captured by covariates “Owner-occupied housing units” and “Total occupied units”. Racial diversity also appears to play a role in cluster formation – in particular, the covariates related to ethnic groups characterize the Ethnic Enclave clusters. Similarly, covariates related to poverty, for example, “Persons below the poverty line”, lack of education (“No high school diploma”), and lack of health insurance (“No health insurance”), determine the Economically Disadvantaged clusters. Covariates related to high mobility, for example, “Different housing unit in prior year”, high education (“College or higher degree holders”), marital status (“Single person households”), and nonspousal occupants (“Unmarried couples”), characterize the Single Unattached Mobiles cluster.
@@
Some of the covariates in Figure 7[right] can also be interpreted in terms of the mindset solutions in Kulzick et al. (2019). In a mindset solution, survey respondents are categorized using factor analysis into six mindsets based on their predicted self-response patterns and demographic characteristics. These mindsets are determined by demographic characteristics such as age, race, income, home-ownership, presence of children in the household, marital status, internet use, English proficiency, and country of birth. The six mindsets are: Eager Engagers, Fence Sitters, Individuals with Confidentiality, Head Nodders, Wary Skeptics and Disconnected Doubters. The covariates can be also interpreted in terms of eight segmented tracts: Responsive Suburbia, Main Street Middle, Country Roads, Downtown Dynamic, Student and Military Communities, Sparse Spaces, Multicultural Mosaic, and Rural Delta and Urban Enclaves – see Kulzick et al. (2019) for additional details.
@@ 5.4 Conclusions
We propose a framework to obtain models with good predictive accuracy while simultaneously delivering parsimonious and interpretable statistical models. We hope that our framework will help practitioners identify some key factors that influence survey response rates. Insights gathered from our models may be used for targeted follow-up surveys and resource allocation for advertisements. We hope that our models will help inform the goal of achieving improved census coverage in a timely fashion while optimizing operational risks and costs. Our framework can complement and potentially offer an alternative to the LRS metric currently used in the US Census Bureau ROAM application.
@@ Acknowledgements
The authors would like to thank Nancy Bates, Tommy Wright and Joanna Fane Lineback from the U.S. Census Bureau for helpful discussions. Thanks also to other researchers from the Centers for Statistical Research and Methodology and Behavioral Science Methods (U.S Census Bureau) for helpful feedback in earlier versions of this work. Any opinions and conclusions expressed herein are those of the authors and do not necessarily represent the views of the U.S. Census Bureau. This research was supported in part, by grants from the Office of Naval Research: N000141812298, N000142112841 and N000142212665 and the National Science Foundation: NSF-IIS-1718258 awarded to Rahul Mazumder.
[…]
Supplementary Material: Predicting Census Survey Response Rates with Parsimonious Additive Models and Structured Interactions
by Shibal Ibrahim, Peter Radchenko, Emanuel Ben-David, and Rahul Mazumder
@@
We study how our model performs when the error distribution is (i) skewed, (ii) heteroskedastic. For (i), we consider a log-normal distribution for the error: $\epsilon\sim\mathcal{LN}(-1,0.2546)$. For (ii), we consider a normal distribution for the error: $\epsilon\sim\mathcal{N}(0,0.5092g_{2}(x_{5}))$, where the standard deviation is a function of the 5-th covariate in $\mathbf{x}$. We compare average ISE for ELAAN-I with EBM and GAMI-Net in these scenarios in Table S2. We can observe that ELAAN-I consistently has favorable performance over both EBM and GAMI-Net even when the error distributions are not normally distributed. As expected, we observe a small drop in performance in both skewed and heteroskedastic settings compared to the normal distribution setting. We do not observe any degradation in support recovery metric in Table S3 under skewed setting; however, we do observe a drop in support recovery F1-scores in the heteroskedastic setting. For example, the F1-scores for main and interaction effects for ELAAN-I are $96.4\%$ and $81.7\%$ in the heteroskedastic setting for $N_{train}=400$ compared to $97.4\%$ and $87.3\%$ in the normal distribution setting.
@@
For EBM, we tune the number of interactions in the range $[1,100]$. For GAMI-Net, we set the number of interactions in to $500$. For GAMI-Net, we used $200$ batch-size, $0.001$ learning rate, $500$ epochs (for each stage). Loss threshold was set to $0.001$, which controls how many effects are selected after training and pruning. For GAMI-Net, we considered 2-layered NNs with 10 neurons per layer for each effect as suggested by the referee. The default basis with 5-layered NN with 40-neurons per layer was much slower.